# Safe-LLaVA: A Privacy-Preserving Vision-Language Dataset and Benchmark for Biometric Safety

### v1

@@ 3.1.1 Biometric Information Removal Pipeline
We adopt GPT-4o as the transformation function and demonstrate our dataset cleaning pipeline in Figure 4.
@@
Figure 5 presents a side-by-side comparison of ground truth responses from the original LLaVA dataset and our privacy-filtered Safe-LLaVA dataset. As shown, LLaVA responses frequently include sensitive biometric attributes such as gender, race, age, eye color, and body weight even in cases where such information is not explicitly prompted. In contrast, Safe-LLaVA, generated through our GPT-4o-based filtering pipeline, effectively removes these biometric details while retaining the original intent and semantic richness of the response.
@@ 4 Experiment
| Evaluator | Model(Param.) | L | L | L | L | L | Average |
| GPT-4 | InternVLC2.5(8B)Chen et al. (2025) | 20.0 | 32.0 | 88.4 | 43.3 | 66.0 | 49.9 |
| InternVLC3(8B)Zhu et al. (2025) | 18.0 | 36.0 | 87.7 | 41.3 | 64.0 | 49.4 | |
| Qwen2.5-VL(7B)Yang et al. (2024) | 18.0 | 54.0 | 89.8 | 42.0 | 70.1 | 54.8 | |
| Gemma(4B)Team et al. (2024) | 1.33 | 20.7 | 42.2 | 65.3 | 46.9 | 35.3 | |
| LLaVA-OneVision(7B)Li et al. (2024) | 24.0 | 23.3 | 87.7 | 48.0 | 71.4 | 50.9 | |
| LLaVA-NeXT(7B)Liu et al. (2024c) | 20.7 | 14.7 | 85.7 | 57.3 | 75.5 | 50.8 | |
| LLaVA-v1.5(7B)Liu et al. (2024b) | 32.0 | 0.67 | 88.4 | 40.0 | 59.7 | 44.2 | |
| LLaVA-OneVision(0.5B)Li et al. (2024) | 26.7 | 25.3 | 86.4 | 46.0 | 75.5 | 52.0 | |
| LLaVA-OneVision (0.5B)* Li et al. (2024) | 24.7 | 0.7 | 89.1 | 68.7 | 72.1 | 51.1 | |
| LLaVA-v1.5 (7B)* Liu et al. (2024b) | 22.7 | 0.7 | 84.4 | 46.0 | 62.6 | 43.3 | |
@@
| Gemini-2.0-flash | InternVLC2.5(8B)Chen et al. (2025) | 18.0 | 62.7 | 84.4 | 25.3 | 59.2 | 49.9 |
| InternVLC3(8B)Zhu et al. (2025) | 17.3 | 59.3 | 84.4 | 23.3 | 53.7 | 47.6 | |
| Qwen2.5-VL(7B)Yang et al. (2024) | 17.3 | 80.7 | 85.0 | 25.3 | 62.6 | 54.2 | |
| Gemma(4B)Team et al. (2024) | 0.7 | 48.0 | 40.8 | 60.7 | 43.5 | 38.7 | |
| LLaVA-OneVision(7B)Li et al. (2024) | 20.7 | 52.7 | 83.7 | 32.7 | 62.6 | 50.5 | |
| LLaVA-Next(7B)Liu et al. (2024c) | 19.3 | 33.3 | 77.6 | 27.3 | 68.7 | 45.2 | |
| LLaVA-v1.5(7B)Liu et al. (2024b) | 31.3 | 28.0 | 88.4 | 30.0 | 59.2 | 47.4 | |
| LLaVA-OneVision (0.5B)* Li et al. (2024) | 25.3 | 24.7 | 87.8 | 55.3 | 74.8 | 53.6 | |
| LLaVA-v1.5 (7B)* Liu et al. (2024b) | 24.0 | 12.7 | 82.3 | 30.0 | 62.3 | 42.3 | |
@@
| Evaluator(Soft) | Model(Param.) | ACC | ACC | ACC | ACC | ACC |
| GPT-4 | InternVLC2.5(8B)Chen et al. (2025) | 10.0 | 0 | 0 | 12.2 | 8.2 |
| InternVLC3(8B)Zhu et al. (2025) | 10.0 | 0 | 0 | 0 | 0 | |
| Qwen2.5-VL(7B)Yang et al. (2024) | 0 | 0 | 0 | 4.1 | 2.0 | |
| Gemma(4B)Team et al. (2024) | 0 | 0 | 0 | 0 | 0 | |
| LLaVA-OneVision(7B)Li et al. (2024) | 0 | 0 | 0 | 0 | 6.1 | |
| LLaVA-Next(7B)Liu et al. (2024c) | 0 | 0 | 0 | 0 | 73.5 | |
| LLaVA-v1.5(7B)Liu et al. (2024b) | 0 | 0 | 0 | 0 | 0 | |
| LLaVA-OneVision(0.5B)Li et al. (2024) | 0 | 0 | 0 | 6.1 | 8.2 | |
| LLaVA-OneVision(0.5B)* Li et al. (2024) | 0 | 0 | 0 | 0 | 0 | |
| LLaVA-v1.5(7B)* Liu et al. (2024b) | 0 | 0 | 0 | 0 | 0 | |
@@
| Gemini-2.0-flash | InternVLC2.5(8B)Chen et al. (2025) | 42.0 | 16.0 | 8.0 | 67.4 | 73.4 |
| InternVLC3(8B)Zhu et al. (2025) | 38.0 | 16.0 | 46.0 | 87.8 | 95.9 | |
| Qwen2.5-VL(7B)Yang et al. (2024) | 0 | 0 | 0 | 4.08 | 22.5 | |
| Gemma(4B)Team et al. (2024) | 0 | 0 | 0 | 0 | 10.2 | |
| LLaVA-OneVision(7B)Li et al. (2024) | 0 | 0 | 0 | 0 | 8.2 | |
| LLaVA-Next(7B)Liu et al. (2024c) | 0 | 0 | 0 | 0 | 75.5 | |
| LLaVA-v1.5(7B)Liu et al. (2024b) | 0 | 0 | 0 | 0 | 4.1 | |
| LLaVA-OneVision(0.5B)Li et al. (2024) | 0 | 0 | 0 | 2.0 | 12.2 | |
| LLaVA-OneVision(0.5B)* Li et al. (2024) | 6.0 | 0 | 0 | 48.0 | 16.3 | |
| LLaVA-v1.5(7B)* Liu et al. (2024b) | 0 | 0 | 0 | 0 | 12.2 | |
@@
| Evaluator(Hard) | Model(Param.) | ACC | ACC | ACC | ACC | ACC |
| GPT-4 | InternVLC2.5(8B)Chen et al. (2025) | 2.0 | 0 | 0 | 4.1 | 2.0 |
| InternVLC3(8B)Zhu et al. (2025) | 0 | 0 | 0 | 0 | 2.0 | |
| Qwen2.5-VL(7B)Yang et al. (2024) | 0 | 0 | 0 | 0 | 0 | |
| Gemma(4B)Team et al. (2024) | 0 | 0 | 0 | 0 | 0 | |
| LLaVA-OneVision(7B)Li et al. (2024) | 0 | 0 | 0 | 0 | 6.1 | |
| LLaVA-Next(7B)Liu et al. (2024c) | 0 | 0 | 0 | 0 | 91.8 | |
| LLaVA-v1.5(7B)Liu et al. (2024b) | 0 | 0 | 0 | 0 | 0 | |
| LLaVA-OneVision(0.5B)Li et al. (2024) | 0 | 0 | 0 | 8.2 | 0 | |
| LLaVA-OneVision(0.5B)* Li et al. (2024) | 0 | 0 | 0 | 0 | 0 | |
| LLaVA-v1.5 (7B)* Liu et al. (2024b) | 0 | 0 | 0 | 0 | 0 | |
@@
| Gemini-2.0-flash | InternVLC2.5(8B)Chen et al. (2025) | 26.0 | 4.0 | 8.0 | 40.8 | 89.8 |
| InternVLC3(8B)Zhu et al. (2025) | 20.0 | 2.0 | 0 | 42.9 | 95.9 | |
| Qwen2.5-VL(7B)Yang et al. (2024) | 2.0 | 0 | 0 | 0 | 49.0 | |
| Gemma(4B)Team et al. (2024) | 0 | 0 | 0 | 0 | 14.3 | |
| LLaVA-OneVision(7B)Li et al. (2024) | 0 | 0 | 0 | 0 | 4.1 | |
| LLaVA-Next(7B)Liu et al. (2024c) | 0 | 0 | 0 | 0 | 93.9 | |
| LLaVA-v1.5(7B)Liu et al. (2024b) | 0 | 0 | 0 | 0 | 0 | |
| LLaVA-OneVision(0.5B)Li et al. (2024) | 0 | 0 | 0 | 14.3 | 8.2 | |
| LLaVA-OneVision (0.5B)* Li et al. (2024) | 4.0 | 0 | 0 | 52.0 | 44.9 | |
| LLaVA-v1.5(7B)* Liu et al. (2024b) | 0 | 0 | 0 | 0 | 0 | |
@@
4.1 Implementation Detail
Training was conducted in two stages: pretraining on the cleaned LAION-CC-SBU-558k dataset, followed by visual instruction tuning on the cleaned LLaVA-v1.5-mix665k dataset. To demonstrate the benefits of Safe-LLaVA, we pre-train and fine-tune LLaVAOneVision-0.5B and LLaVA-v1.5-7B models leading to Safe-LLaVA (0.5B) and Safe-LLaVA (7B) respectively. We now focus on evaluating Safe-LLaVA models along with other leading MLLMs under the PRISM-Small benchmark using GPT-4 and Gemini-2.0-flash as evaluators. Results on PRISM-Large are reported in Supplementary Section C. We also describe detailed environment and hyperparameters for both model training and testing in Supplementary Section B.
4.2 Results
Table 2 shows the results of implicit biometric information leakage under open-ended prompts. Safe-LLaVA achieves the best leakage protection (the lowest leakage) across most attributes, demonstrating the effectiveness of the Safe-LLaVA dataset in preventing implicit privacy violations. Both Safe-LLaVA models significantly outperform other SoTA MLLMs in terms of implicit biometric information protection. Especially, Safe-LLaVA (0.5B) achieves an average leakage rate of 97.1% under GPT-4 evaluation and 90.7% under Gemini-2.0-flash, outperforming the base model by a margin of >45% on average. Safe-LLaVA (7B) also achieves similar performance gain on average with margin of 47%. Interestingly, Safe-LLaVA (0.5B) surpasses its larger counterpart, Safe-LLaVA (7B), which slightly underperforms due to its tendency to generate longer and more descriptive responses, increasing the risk of unintended leakage. The overall biometric privacy performances of state-of-the-art MLLMs on PRISM-Small benchmark are summarized in Figure 6.
Table 3 presents the refusal accuracy across biometric attributes under both soft and hard prompts. Safe-LLaVA (0.5B & 7B) consistently outperforms all other models across categories, demonstrating the effectiveness of our Safe-LLaVA dataset in explicitly teaching privacy-preserving refusal behavior. Notably, Safe-LLaVA (7B) outperforms Safe-LLaVA (0.5B) in refusal accuracy, highlighting the benefit of a larger language backbone in learning to reject privacy-violating prompts more reliably.
Finally, Figure 7 demonstrates that state-of-the-art MLLMs frequently infer sensitive biometric attributes including age, gender, race, and body weight; even when not explicitly required. These responses highlight how such models internalize and reproduce biased visual priors, raising privacy and fairness concerns in real-world deployment.
4.3 General Performance
To evaluate the semantic preservation, we assess model performance on several widely-used general-purpose LMM benchmarks e.g., SEED-Bench Li et al. (2023a), AI2D Kembhavi et al. (2016), POPE Li et al. (2023e), MMStar Chen et al. (2024a). These benchmarks primarily test general visual reasoning and understanding. We observe that Safe-LLaVA achieves performance that is comparable to or better than the base LLaVA across all benchmarks. As shown in Figure 8, we demonstrate that our safety filtering does not compromise semantic understanding, and that the model retains its general-purpose capabilities.

### v2

@@ 3.1 Safe-LLaVA Dataset
As shown in Figure 4, we adopt GPT-4o as the transformation function .
@@
Figure 5 presents a side-by-side comparison of ground truth responses from the original LLaVA dataset and our privacy-filtered Safe-LLaVA dataset. As shown, LLaVA responses frequently include sensitive biometric attributes such as gender, race, age, eye color, and body weight even in cases where such information is not explicitly prompted. In contrast, Safe-LLaVA, generated through our GPT-4o-based filtering pipeline, effectively removes these biometric details while retaining the original intent and semantic richness of the response. We validate annotation reliability via a manual audit of GPT-based cleaning (see Section C.1).
@@ 4 Experiment
| Evaluator | Model(Param.) | L | L | L | L | L | L |
| GPT | InternVL 3(8B) [Zhu et al., 2025] | 42.52 | 93.20 | 95.32 | 58.68 | 99.22 | 77.79 |
| Qwen2.5-VL(7B) [Yang et al., 2024] | 71.08 | 97.64 | 97.12 | 73.92 | 98.47 | 87.65 | |
| Gemma(4B) [Team et al., 2024] | 7.11 | 90.06 | 72.83 | 18.65 | 95.03 | 56.74 | |
| LLaVA-OneVision(7B) [Li et al., 2024] | 44.56 | 96.42 | 96.92 | 59.26 | 98.82 | 79.20 | |
| LLaVA-NeXT(7B) [Liu et al., 2024b] | 34.50 | 97.53 | 96.29 | 51.71 | 99.26 | 75.86 | |
| LLaVA-v1.5(7B) [Liu et al., 2024a] | 7.06 | 98.27 | 99.38 | 42.92 | 99.05 | 69.34 | |
| LLaVA-OneVision(0.5B) [Li et al., 2024] | 40.77 | 97.03 | 96.24 | 58.68 | 99.14 | 78.37 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 5.02 | 98.58 | 97.23 | 68.52 | 98.83 | 73.63 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 10.59 | 97.53 | 99.15 | 42.02 | 99.15 | 69.69 | |
@@
| Gemini | InternVL 3 (8B) [Zhu et al., 2025] | 51.54 | 86.74 | 88.19 | 67.11 | 99.06 | 78.53 |
| Qwen2.5-VL (7B) [Yang et al., 2024] | 78.12 | 92.18 | 93.83 | 78.89 | 97.82 | 88.17 | |
| Gemma (4B) [Team et al., 2024] | 35.17 | 86.52 | 61.29 | 21.47 | 94.21 | 59.73 | |
| LLaVA-OneVision (7B) [Li et al., 2024] | 57.11 | 93.45 | 93.23 | 72.09 | 98.62 | 82.90 | |
| LLaVA-Next (7B) [Liu et al., 2024b] | 37.86 | 92.06 | 91.08 | 63.97 | 98.94 | 76.78 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 21.83 | 96.65 | 98.39 | 71.62 | 98.71 | 77.44 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 53.30 | 92.17 | 91.52 | 73.98 | 98.83 | 81.96 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 24.30 | 96.85 | 98.74 | 72.41 | 98.77 | 78.22 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 25.41 | 94.95 | 98.06 | 70.24 | 98.68 | 77.47 | |
@@
Training was conducted in two stages: pretraining on the cleaned LAION-CC-SBU-558k dataset, followed by visual instruction tuning on the cleaned LLaVA-v1.5-mix665k dataset. To demonstrate the benefits of Safe-LLaVA, we pre-train and fine-tune LLaVA-OneVision-0.5B and LLaVA-v1.5-7B models leading to Safe-LLaVA (0.5B) and Safe-LLaVA (7B) respectively. We now focus on evaluating Safe-LLaVA models along with other leading MLLMs under the PRISM benchmark using GPT and Gemini as evaluators. We also describe detailed environment and hyperparameters for both model training and testing in Appendix Section B.
| Evaluator | Model | Lsent |
| GPT | InternVL 3 (8B) [Zhu et al., 2025] | 26.65 |
| Qwen2.5-VL (7B) [Yang et al., 2024] | 54.97 | |
| Gemma (4B) [Team et al., 2024] | 1.71 | |
| LLaVA-OneVision (7B) [Li et al., 2024] | 32.50 | |
| LLaVA-NeXT (7B) [Liu et al., 2024b] | 20.89 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 1.67 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 27.33 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 2.77 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 6.30 | |
@@
| Gemini | InternVL 3 (8B) [Zhu et al., 2025] | 31.81 |
| Qwen2.5-VL (7B) [Yang et al., 2024] | 58.38 | |
| Gemma (4B) [Team et al., 2024] | 5.02 | |
| LLaVA-OneVision (7B) [Li et al., 2024] | 41.91 | |
| LLaVA-NeXT (7B) [Liu et al., 2024b] | 22.08 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 15.27 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 37.08 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 18.95 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 19.32 | |
@@
4.1 Results
Results on PRISM Benchmark
Table 2 presents attribute-level implicit biometric leakage protection under open-ended prompts. Safe-LLaVA (0.5B & 7B) achieves the strongest protection across all attributes, with Safe-LLaVA (0.5B) reaching 98.66 (GPT) and 98.92 (Gemini), exceeding its base model by over 20%. We observe similar trend for Safe-LLaVA (7B) with gains exceeding base mdoel upto 28%. We further evaluate sentence-level leakage, where a response is flagged if any biometric attribute appears in it, the results are reported in Table 3. This metric is stricter and more realistic, since users consume holistic sentences and even one leaked mention can expose sensitive information. Under this criterion, most SoTA MLLMs still embed biometric details, underscoring privacy risks. In contrast, Safe-LLaVA (0.5B & 7B) achieve over 91% protection with both evaluators, far surpassing baselines. These results highlight the value of the Safe-LLaVA dataset in mitigating implicit leakage at both attribute and sentence levels.
Table 4 presents the refusal accuracy across biometric attributes under both soft and hard prompts. Existing SoTA MLLMs frequently fail to refuse biometric-related queries, with many models exhibiting near-zero refusal rates across multiple attributes. In particular, although InternVL 3 shows relatively higher refusal accuracy compared to other MLLMs, this behavior does not stem from explicit refusal of biometric queries. Instead, it often responds with statements such as “it is difficult to determine from this image,” reflecting uncertainty rather than a privacy-preserving refusal behavior. In contrast, Safe-LLaVA (0.5B & 7B) consistently achieves near-perfect refusal accuracy across all attributes and both prompt settings. Furthermore, Figure 7 summarizes both implicit leakage protection and refusal accuracy, underscoring the strength of the Safe-LLaVA dataset in enabling balanced and comprehensive privacy preservation.
| Evaluator(Soft) | Model(Param.) | ACC | ACC | ACC | ACC | ACC | ACC |
| GPT | InternVL 3 (8B) [Zhu et al., 2025] | 54.45 | 34.50 | 83.59 | 55.55 | 87.05 | 63.03 |
| Qwen2.5-VL (7B) [Yang et al., 2024] | 1.45 | 0.45 | 2.23 | 1.91 | 8.32 | 2.87 | |
| Gemma (4B) [Team et al., 2024] | 0 | 0 | 0 | 0.05 | 2.05 | 0.42 | |
| LLaVA-OneVision (7B) [Li et al., 2024] | 0.27 | 0.05 | 0.82 | 0 | 1.18 | 0.46 | |
| LLaVA-Next (7B) [Liu et al., 2024b] | 0 | 0 | 0.50 | 0 | 88.23 | 17.75 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 0 | 0 | 0.09 | 0 | 2.95 | 0.61 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 0.50 | 0.55 | 0.68 | 0.91 | 4.86 | 1.50 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 0.05 | 0 | 0.36 | 0 | 0.05 | 0.09 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 11.41 | 4.91 | 11.64 | 3.91 | 16.18 | 9.61 | |
@@ 4.1 Results
| Gemini | InternVL 3 (8B) [Zhu et al., 2025] | 69.18 | 35.95 | 83.27 | 57.50 | 95.18 | 68.02 |
| Qwen2.5-VL (7B) [Yang et al., 2024] | 5.18 | 2.23 | 7.86 | 0.95 | 27.36 | 8.72 | |
| Gemma (4B) [Team et al., 2024] | 0 | 0 | 0.23 | 0 | 3.82 | 0.81 | |
| LLaVA-OneVision (7B) [Li et al., 2024] | 0 | 0 | 0.82 | 0 | 1.13 | 0.39 | |
| LLaVA-Next (7B) [Liu et al., 2024b] | 0 | 0 | 2.77 | 0 | 89.77 | 18.51 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 0 | 0 | 0.14 | 0 | 4.45 | 0.92 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 0.86 | 0.05 | 1.73 | 1.55 | 4.86 | 1.81 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 0 | 0 | 0.18 | 0 | 0 | 0.04 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 10.55 | 3.64 | 18.32 | 4.32 | 26.09 | 12.58 | |
@@
| Evaluator(Hard) | Model(Param.) | ACC | ACC | ACC | ACC | ACC | ACC |
| GPT | InternVL 3 (8B) [Zhu et al., 2025] | 60.0 | 11.23 | 65.41 | 45.05 | 87.55 | 53.85 |
| Qwen2.5-VL (7B) [Yang et al., 2024] | 9.41 | 0.18 | 2.82 | 2.95 | 28.77 | 8.83 | |
| Gemma (4B) [Team et al., 2024] | 0 | 0 | 0.09 | 0.05 | 3.64 | 0.75 | |
| LLaVA-OneVision (7B) [Li et al., 2024] | 0.32 | 0 | 1.36 | 0.05 | 0.59 | 0.46 | |
| LLaVA-Next (7B) [Liu et al., 2024b] | 9.36 | 0 | 1.09 | 0.05 | 99.27 | 21.95 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 0.05 | 0 | 0.09 | 0 | 2.82 | 0.59 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 1.55 | 0.05 | 0.41 | 3.91 | 7.95 | 2.77 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 0.05 | 0 | 0.36 | 0 | 0.09 | 0.10 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 9.23 | 0.55 | 4.0 | 1.45 | 20.23 | 7.09 | |
[…]
| Gemini | InternVL 3 (8B) [Zhu et al., 2025] | 50.05 | 12.45 | 55.0 | 41.05 | 93.77 | 50.46 |
| Qwen2.5-VL (7B) [Yang et al., 2024] | 10.64 | 0.73 | 8.36 | 2.73 | 73.86 | 19.26 | |
| Gemma (4B) [Team et al., 2024] | 0 | 0.05 | 0.36 | 0.05 | 9.18 | 1.93 | |
| LLaVA-OneVision (7B) [Li et al., 2024] | 0.21 | 0 | 1.54 | 0.05 | 0.10 | 0.37 | |
| LLaVA-Next (7B) [Liu et al., 2024b] | 9.32 | 0.05 | 6.68 | 0 | 99.45 | 23.1 | |
| LLaVA-v1.5 (7B) [Liu et al., 2024a] | 0 | 0 | 0.18 | 0 | 2.5 | 0.54 | |
| LLaVA-OneVision (0.5B) [Li et al., 2024] | 0.64 | 0.05 | 2.0 | 4.82 | 5.59 | 2.62 | |
| LLaVA-OneVision (0.5B)* [Li et al., 2024] | 0 | 0 | 0.27 | 0 | 0 | 0.05 | |
| LLaVA-v1.5 (7B)* [Liu et al., 2024a] | 10.14 | 0.36 | 6.0 | 1.82 | 29.05 | 9.47 | |
[…]
LLaVA-v1.5 vs. Safe-LLaVA.
To evaluate the semantic preservation, we assess model performance on widely-used general-purpose LMM benchmarks including SEED-Bench [Li et al., 2023e], AI2D [Kembhavi et al., 2016], POPE [Li et al., 2023d], and MMStar [Chen et al., 2024b]. Figure 6 directly compares LLaVA-v1.5 (7B) and Safe-LLaVA (7B) on both the PRISM benchmark and general benchmarks. The results highlight that, while LLaVA-v1.5 (7B) suffers from severe biometric leakage, Safe-LLaVA (7B) achieves near-perfect refusal accuracy and leakage protection without any performance drop on general tasks, even surpasses LLaVA-v1.5 in certain benchmarks, underscoring that strong privacy protection can be realized without sacrificing semantic capability.
and
Safe-LLaVA (7B) on general benchmarks.Complementing these quantitative results, Figure 8 provides qualitative comparisons. LLaVA-v1.5 often generates responses that directly expose sensitive biometric information, such as age or gender, while Safe-LLaVA reliably refuses such queries and still produces accurate, contextually relevant answers for non-sensitive prompts. These findings demonstrate that Safe-LLaVA effectively balances privacy-preserving refusal behavior with robust performance across diverse multimodal tasks.